OpenStudy (anonymous):

solve for 2cos2x-(square root 3)=0, please explain why

OpenStudy (adamaero):

cos2x = many different things

OpenStudy (adamaero):

look at your double angle formulas

OpenStudy (anonymous):

2*cos(2*x) = sqrt(3) cos(2*x) = sqrt(3)/2 2*x = 30º x = 15º